OpenStudy (anonymous):

y^6+x^3=y^2+10x ,normal at(0,1) at the given point,find the slope of the curve,the line that is tangent to the curve,or the line that is normal to the curve,as requested.

OpenStudy (amistre64):

the derivative defines the slope ... the equation of a line is from algebra classes so should already be known a normal is a perp line, a line with a perpendicular slope

OpenStudy (anonymous):

this is the problem i could not find it.

OpenStudy (amistre64):

assuming that y is a function of x .... do an implicit y^6+x^3=y^2+10x 6y^5 y' +3x^2 x' = 2y y' + 10x' with respect to x, x'=dx/dx = 1 6y^5 y' +3x^2 = 2y y' + 10, we know (x=0, y=1) 6y'= 2y' + 10 , solve for y', to determine the slope at (0,1)

OpenStudy (amistre64):

now you have a slope, and a point ( a y intercept at that) ... make the equation.

OpenStudy (anonymous):

y=10/4,y=5/2

OpenStudy (anonymous):

y-1=5/2(x-0)=5/2x+1

OpenStudy (amistre64):

good, now what property do perp lines have? something about the slope ...

OpenStudy (anonymous):

It is showing me answer supposed to be y=-2/5x+1

OpenStudy (amistre64):

perp lines have negative reciprocal slopes, yes .. in other words, flip and negate

OpenStudy (amistre64):

slope 5/2, for the tangent slope -2/5, for the normal

OpenStudy (anonymous):

It is very confusing to me.I need to surgery to shift my brain for this problems.:)

OpenStudy (anonymous):

Thank you for your help.

OpenStudy (amistre64):

:) since the problem is asking for the normal line at(0,1) slope of the tangent is y' slope of the normal is -1/y'

OpenStudy (anonymous):

That' great,I need to explain everything how I did it,This is really help.

OpenStudy (amistre64):

good luck

OpenStudy (anonymous):

Tnx.